OpenStudy (anonymous):

Subtract. Express your answer in simplest terms. 4/5 - 1/3

OpenStudy (anonymous):

Exactly what we did in the last one, the only difference is that you will subtract once you have them converted into their common denominator forms.

OpenStudy (anonymous):

well the LCD is 15

OpenStudy (anonymous):

Riight

OpenStudy (anonymous):

4/5 x 3/3 = 12/15

OpenStudy (anonymous):

1/3 x 5/5 = 5/15

OpenStudy (anonymous):

7/15

OpenStudy (anonymous):

Correct, good job

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

no problem